Find the probability of this trial, x = 8, n = 11, p = 0.9, where x is successes in n trials.
Answer by mathmate(429)

Question:
Find the probability of this trial, x = 8, n = 11, p = 0.9, where x is successes in n trials.

Solution:
Details are not given for the question, but from the look of it, I will assume that trials are independent of each other, probability is known and remain constant, trials are Bernoulli trials and the number of trials is known and remains constant.

With these conditions, the binomial distribution may be prescribed, with
P(x)=C(n,x)p^(x)(1-p)^(n-x)
and
C(n,x)=number of combinations of x items taken from n=n!/(x!(n-x)!)

Putting in numbers,
n=11, p=0.9, x=8
P(8)=C(11,8)0.9^8 (1-0.9)^(11-8)
=165*0.430467*0.001
=0.071027